OpenStudy (anonymous):

Please help!! I dont know where to start If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x - 5y + z = 12 x + 6y - z = -4 2x - 5y - 2z = -8

OpenStudy (tkhunny):

Have you considered solving one equation for a variable of your choice and substituting the result into the other two equations?

OpenStudy (anonymous):

Read the instructions, then follow along.... Isolate x in the second equation gives\[x =- 6y+ z -4\]then substitute....\[3x - 5y + z = 12 \implies 3(- 6y+ z -4) - 5y + z = 12\]\[2x - 5y - 2z = -8 \implies 2(- 6y+ z -4) - 5y - 2z = -8\]Simplify, and you have your system.